# Prediction of causal genes at GWAS loci with pleiotropic gene regulatory effects using sets of correlated instrumental variables

### v1

@@ Results
We observe that the GMM estimator is identical to the two-stage least squares estimator of [20] and [10].
@@
We further tested the effect of weak instruments, that is, instruments with weak effects on the exposures (see Methods for details). As expected, weak instruments increase the variance in the causal effect estimates, more strongly at small sample sizes ($n=500-1,000$) (Fig. 2D,E). In general, the increase in variance due to weak instruments is not greater in the MVMR setting than in standard, univariate MR (Supp. Fig. S1B,D).
Finally, we tested the effect of missing exposures in the model, by applying univariate MR to data simulated from a model with two correlated exposures. Unsurprisingly, in this case the effect estimates are highly biased (Fig. 2G,H). In contrast, when only true instruments (causal variants) are missing from the estimation in the over-determined case, effect estimates remain unbiased with only a mild increase in variance (Supp. Fig. S1G).
@@
In our next set of experiments, we compared the LS and GMM estimators in both exactly determined and overdetermined systems, using more realistic correlation matrices taken from real LD values and eQTL and GWAS summary statistics at real loci in the human genome (Fig. 3; see Methods for details). While both estimators are guaranteed to be consistent, the GMM estimator is the estimator with theoretically lowest variance (see Methods). However, across thousands of simulations with multiple sample sizes and for models of varying degree of overdetermination, we did not observe any differences in variance between the LS and GMM estimators (Fig. 3B,C).
We also compared the LS and GMM estimators to estimators from TWMR [10], TwoSampleMR (function mv_multiple for multivariable MR analysis) [31] and the MVMR R package [30]. In theory, the two-stage least squares estimators used in the TWMR and MVMR R packages should be identical to our GMM estimator. This is indeed what we observe for the MVMR R package (Fig. 3D,E,G,H). However, for TWMR we observe systematically that its estimate is not consistent, that is, the average estimated effect over 1,000 independent simulations with the same parameter settings is not equal to the true causal effect (Fig. 3D,E,G,H). Inspection of the TWMR source code revealed the presence of a “regularization” term that was hard-coded to a specific value, related to the sample size of the use-case from their paper [10] (see Methods for details). After removing this term, the TWMR estimates again coincided with our GMM estimator (data not shown).
Finally, in line with the previous set of experiments, we did not observe any reduction in variance of the GMM or MVMR R package estimators compared to the LS estimator, although the GMM estimator exhibits faster convergence to the true causal effect. However, we also observed that the LS estimator demonstrated greater robustness (less increase in variance) to high LD values, which are typical of SNPs colocated in the same genomic locus.
@@
To test our method on real-world data, we predicted causal genes at genome-wide significant loci for coronary artery disease (CAD) from the CARDIOGRAMC4D meta-analysis study [32] using eQTL data from seven vascular and metabolic tissues from the STARNET study [26] (see Methods for details). Results are presented for the least squares estimator; the GMM estimates are predominantly consistent and are provided in the Supplementary Information.
@@
We confirmed the validity of this threshold using seven genome-wide significant CAD loci which had an effect on expression of a single candidate gene in at least two tissues in the STARNET data, that is, loci where we can plausibly hypothesize that the true causal gene, but not the tissue, is known. We used standard, univariate MR on a tissue-by-tissue basis and found that in all seven cases, the candidate gene had a predicted absolute causal effect size greater than 0.1 in at least one tissue (Fig. 4B, Supplementary Information). These genes included CDKN2B, located in one of the most robust genetic markers for type 2 diabetes, CAD, and myocardial infaction [35]. GWAS SNPs in the CDKN2B locus were associated with expression of CDKN2B in aorta (AOR), subcutaneous fat (SF), and blood, with predicted causal effects $c=-0.75$, $-0.55$, and $0.5$, respectively.
@@
Having found a reasonable threshold to define causal genes for CAD, we analyzed 12 genome-wide significant CAD GWAS loci which had an effect on expression of multiple nearby genes in one or more tissues. Similar to [10], we applied MVMR on a tissue-by-tissue basis using all genes with cis-eQTL associations in a locus in a given tissue as potential exposures in a causal diagram (cf. Fig. 1B,E). We note that at only two of the 12 loci, it was possible to reduce the SNPs to a set of independent instruments using the LD threshold ($r^{2}\leq 0.1$) used by [10], while still retaining at least as many instruments as exposures. Hence an approach using correlated instrument sets is essential.
@@
In total, 35 candidate causal genes with cis-eQTL associations were located in the 12 CAD GWAS loci with pleiotropic gene regulatory effects, of which 24 had a predicted absolute causal effect size greater than 0.1 in at least one tissue (Fig. 4D, Supplementary Information). To analyze these causal genes more systematically, we compared the predicted causal effects from MVMR against the predicted causal effects from standard MR (Fig. 4E) and regulatory trait concordance (RTC) scores [36] computed by [26] (Fig. 4F), two univariate methods that ignore the pleiotropic gene regulatory effects. We also conducted a replication analysis for loci with cis-associations in matching tissues in both STARNET and GTEx (Fig. 4G, see Methods for details).
@@
A clear example of a CAD GWAS locus where MVMR and MR give contrasting results is a locus centred around chr 6:12,901,441. In STARNET, the locus has cis-associations with the expression of PHACTR1, GFOD1, TBC1D7, RP1-257A7.4, and RP1-257A7.5 in MAM (Fig. 5A). In univariate MR analyses these genes have predicted causal effects ranging from 0.15 for PHACTR1 to 0.31 for GFOD1, and a previous integrative genomics analysis lists all of them as candidate causal genes for this locus [37]. After pruning nearly identical SNPs (LD $r^{2}\geq 0.95$), 12 instruments (with mutual LD ranging from $r^{2}=0.17$ to $r^{2}=0.86$) were available for conducting MVMR, resulting in a predicted causal effect of $0.19$ for PHACTR1, with all other effects below the threshold of $0.1$ in absolute value, suggesting that PHACTR1 is the only causal gene in arterial tissue at this locus.
@@
Another example is the locus centred around chr 15:79,141,784, where functional studies support a causal and proatherogenic role for ADAMTS7 [38]. In STARNET, the locus has cis-associations with the expression of ADAMTS7 in AOR, MAM, and VAF. In both arterial tissues. In all three tissues, the locus also has cis-associations with the expression of CTSH. After pruning nearly identical SNPs (LD $r^{2}\geq 0.86$), 9, 25, and 10 instruments (with mutual LD ranging from $r^{2}=0.03$ to $r^{2}=0.94$) were available for conducting MVMR in AOR, MAM, and VAF respectively. In the arterial tissues, MVMR confirms that ADAMTS7 is the most likely causal gene (predicted causal effects $0.18$ and $0.14$ vs. $0.025$ and $-0.058$ for CTSH in AOR and MAM, respectively; Fig. 5B). Replication in GTEx AOR tissue confirmed these results (predicted causal effects $0.18$ and $0.009$ for ADAMTS7 and CTSH, respectively). By contrast, in VAF, both genes are predicted causal with effects in the opposite direction (predicted causal effects $-0.18$ and $0.25$ for ADAMTS7 and CTSH, respectively), the latter due to opposite eQTL effects of the same variants in VAF vs. AOR and MAM. No eQTL associations with either gene were available in GTEx in adipose tissue to confirm these results. The overall picture is further complicated by the fact that the locus also has cis-associations with the expression of CTSH in blood, SKLM, and SF, and with PSM4 in SKLM. The absence of associations with ADAMTS7 in these tissues implies that the association between the locus and CAD is automatically attributed to the other genes (Fig. 5B). Given the overwhelming functional evidence for ADAMTS7 [38], these results probably reaffirm the importance of having all true exposures included in the model (cf. Fig. 2F-H).
@@
As a final examples we consider the locus centred around chr 2:85,809,989. Previous analyses found a candidate causal SNP in this locus located in the promoter region of MAT2A, but since the SNP was associated to expression of multiple genes in this locus, a causal gene could not be predicted [39]. In STARNET, the locus has cis-associations with the expression of MAT2A in AOR and MAM. In both tissues, the locus also has cis-associations with the expression of GGCX, and both genes have previously been listed as candidate causal genes in arterial tissues [37]. After pruning nearly identical SNPs (LD $r^{2}\geq 0.95$), three instruments (with mutual LD ranging from $r^{2}=0.69$ to $r^{2}=0.89$) were available for conducting MVMR. In both AOR and MAM, MVMR suggests MAT2A, and not GGCX is the causal gene (predicted causal effects $-0.23$ and $-0.33$ vs. $0.02$ and $0.16$ for GGCX in AOR and MAM, respectively; Fig. 5B). The overall picture is again less clear, since the locus also has cis-associations with the expression of GGCX and VAMP8 in liver, and with GGCX alone in SF and VAF. Interestingly, in liver, neither gene is predicted causal (predicted causal effects $-0.08$ and $-0.06$; Fig. 5B), while in the adipose tissues, the entire association between the locus and CAD is automatically attributed to GGCX (predicted causal effects $-0.14$ in both tissues; Fig. 5B).
@@ Discussion
Using the method of path coefficients we showed that the MVMR causal diagram with correlated instruments satisfies the instrumental set condition, a classical result by [24] that guarantees the identifiability of the direct causal effects of all exposures in the model on the outcome variable of interest. Moreover, the effects solve a regression of (univariate) regression coefficients, and the only requirement is that the set of instruments consists of (or tags) at least as many causal variants as the number of candidate genes included in the model. We further showed that the standard two-stage least squares estimator of the causal effects in MVMR is part of a generalized method of moments (GMM) family of finite-sample estimators for the theoretical identification equation, as is a least squares estimator that simply replaces the theoretical univariate regression coefficients by their finite sample estimates, results which again do not depend on the presence or absence of instrument correlations.
Extensive simulations confirmed the validity and usefulness of these theoretical results even at modest sample sizes ($n\gtrsim 500$-$1000$). Most surprisingly perhaps was the finding that the variance in causal effect estimates remained small, even at high correlation values between the instruments. Moreover we saw no benefit in using the computationally more expensive two-stage least squares estimator instead of the simpler least squares estimator – both were unbiased and had identical variance in all our simulations.
Application of our method to predict causal genes for CAD using eQTL data from seven vascular and metabolic tissues in the STARNET study illustrated the importance of being able to apply MVMR with correlated instrument sets. Out of 36 genome-wide siginificant loci with cis-eQTL associations in STARNET, only 17 were associated with a single gene in a single tissue, 7 with a single gene in multiple tissues, and 12 with multiple genes in multiple tissues. A deeper look in the predictions at some of these loci with widespread regulatory effects showed both the strengths and limitations of applying MVMR in this context.
@@
The fact that regulatory pleiotropy can extend over multiple tissues implies that MVMR must be run strictly speaking by including all gene-tissue pairs with cis-eQTL associations in a given GWAS locus as exposures in a single, large causal model. However, even though instrument sets can be highly correlated and the STARNET data contains only seven tissues, it was almost never possible to run such large models, as the requirement on the number of causal variants included in or tagged by the instrument set still stands. To overcome this limitation, we see two ways forward.
First, we should not consider MVMR to be a hypothesis-free method that can be run across all available tissue eQTL data to discover causal genes de novo. Instead prior knowledge of the relevant tissue and plausible candidate genes must be used to limit both the size of the model and the risk that true causal genes and gene-tissue combinations are missing from the model. For instance, integrative genomics analysis pipelines such as the ones developed by [39] and [37] combine multiple eQTL and GWAS datasets with functional genome annotation and literature search to compile ranked lists of the most likely causal genes and tissues for an outcome of interest such as CAD. Using MVMR as an additional step in such pipelines to resolve causal effects at loci with eQTL effect on prioritized genes in prioritized tissues can provide additional causal evidence, as illustrated by the PHACTR1 and ADAMTS7 examples.

### v2

@@ Results
We observe that the GMM estimator is identical to the two-stage least squares estimator of McDaid et al.[20] and Porcu et al. [10].
@@
Fig 2: Causal effect estimation from data simulating instrument correlations, instrument strength, and horizontal pleiotropy. (A) Causal diagram for the simulation of two exposures $X_{1}$ and $X_{2}$ for an outcome $Y$, with two shared instruments $E_{1}$ and $E_{2}$ with pairwise covariance $\alpha$ and matrix of instrument effect sizes $A=[a_{ij}]$. (B) Difference ($\hat{c}-c$) between estimated and true causal effect of $X_{1}$ (true effect size $c=0.2$) on $Y$ across 1,000 independently simulated datasets for a range of sample sizes. (C) Variance explained by the first two PCs for instrument covariances $\alpha=0.01$, $0.3$, $0.7$, $0.9$ and $0.95$. (D,E) Distribution of the difference between estimated and true causal effects for $X_{1}$ (left; true effect size $c=0.2$) and $X_{2}$ (right; true effect size $c=0.6$) across 1,000 independently simulated datasets and a range of sample sizes comparing strong $0.7-1.0$ vs weak $0.001-0.03$instruments; insets show a zoom along y-axis ranges. (F) False causal diagram used for inference where one causal exposure is missing. (G-H) Difference between estimated and true causal effect of $LPA$ (left, true effect size $c_{\textit{LPA}}=-0.05$) and $PLG$ (right, true effect size $c_{\textit{PLG}}=-0.27$) on $Y$ across 1,000 independently simulated datasets with sample size 2000 across a range of effect sizes of the hidden gene SLC22A3 mediating horizontal pleiotropic effects of the instruments on $Y$, showing both the estimates with the correct model (green) and the estimates with the hidden exposure model (orange). See Methods for simulation details.
Principal component analysis (PCA) has been suggested as a method to prune correlated SNPs in cis-MR studies by using a reduced set of uncorrelated linear combinations of SNPs as instruments instead [34]. We used the multivariable principal component generalized method of moments (MVMR PCA, function mr_mvpcgmm from the MendelianRandomization package) [35] in our simple simulation and found that at an instrument of correlation value of 0.5 or more, the MVMR PCA method reported too few (that is, only one) effective degrees of freedom for causal effect estimation, despite the causal effects still being estimable.
@@
We further tested the effect of weak instruments, that is, instruments with weak effects on the exposures (see Methods for details). As expected, weak instruments increase the variance in the causal effect estimates, more strongly at small sample sizes ($n=500-1,000$) (Fig. 2D,E). In general, the increase in variance due to weak instruments is not greater in the MVMR setting than in standard, univariate MR (S1 Fig B,D).
In our next set of experiments, we used more realistic model parameters taken from real LD correlation matrices and eQTL and GWAS summary statistics from real loci in the human genome (see Methods for details). First we tested the effect of horizontal pleiotropy. Using real summary statistics from the LPA-PLG-SLC22A3 locus, we simulated data from a model where all three genes (exposures) share causal instruments and are causal for the outcome $Y$, but we assumed only data for LPA and PLG was available (Fig. 2F, see Methods for details). In other words, we assumed that the instruments have horizontal pleiotropic effects through the unmeasured gene SLC22A3. As expected, horizontal pleiotropy introduces bias in the estimated causal effects, although the bias remains modest (within the standard deviation of the estimates obtained from the complete model) upto a simulated effect size of around $0.15$ of SLC22A3 on the outcome (Fig. 2G,H).
Next, we tested the effect of LD matrix bias, where the LD matrices in the eQTL and GWAS population differ from the reference LD matrix (e.g. from 1000 Genomes Project). For this purpose, we compared GMM estimates from data generated from the reference LD matrix and data generated from a biased LD matrix sampled from a Wishart distribution centred around the reference LD matrix (see Methods for details), using the reference LD matrix $\Sigma_{EE}$ in both GMM estimates (cf. eq. (10)). With a modest degree of LD matrix bias (Wishart distribution with 50 degrees of freedom), the bias in the causal estimates remained small (Fig. 3B,C).
@@
Next, we compared the LS and GMM causal effect estimators in over-determined models (more instruments than exposures). While both estimators are guaranteed to be consistent, the GMM estimator is the estimator with theoretically lowest variance (see Methods). However, across thousands of simulations with multiple sample sizes and for models of varying degree of over-determination, we did not observe any differences in variance between the LS and GMM estimators (Fig. 3B,C).
Fig 3: Causal effect estimation from data simulating LD matrix sampling bias. Each row shows a causal diagram for the simulation of overdetermined systems where the number of causal variants was equal to the number of exposures (first row, three exposures; second row, two exposures) and other variants were only correlated with the exposures and outcome by LD. (A, B, C) Comparison between the GMM estimator using an unbiased (B) and biased (C) LD matrix for the causal effect of simulated SLC22A3, for different LD pruning thresholds (corresponding to models with 3, 4, 5, 6 and 7 instruments). (D, E, F) Comparison of the least-squares, GMM, TWMR, MV multiple (TwoSampleMR), and MVIVW (Mendelian Randomization package) estimators using an unbiased (E) and biased (F) LD matrix for the causal effect of simulated MRAS and ESTY3. All dsitributions plots show the difference between estimated and true causal effects across 1,000 independently simulated datasets from the causal diagram on the correspinding row. See Methods for simulation details.
We also simulated a two-sample setting where eQTL and GWAS LD matrices were sampled independently from two different Wishart distributions centred around the reference LD matrix, with different variance parameters to reflect the difference in sample size between eQTL and GWAS studies, and generated eQTL and GWAS summary statistics independently, with sample sizes representative of real eQTL and GWAS studies (see Methods for details). In the two-sample setting, even the effects estimated using the true LD matrix from the GWAS population are biased due to the LD matrix sampling variation between the eQTL and GWAS studies (Fig. 3E). Unsurprisingly, the bias increased further when the reference LD matrix was used instead of the GWAS study LD matrix in the estimatio n(Fig. 3F). In this setting, all estimators were biased until an eQTL and GWAS sample size combination of 4,000/140,000 was reached, which corresponds to the upper end of currently available sample sizes (Fig. 3F); TWMR remained biased at all sample sizes, due to a hard-coded regularization factor (see Methods for details). We note that the causal effect bias in two-sample settings is not specific to MVMR or the use of correlated instruments, but can already be observed in univariate MR with a single instrument (S5 Fig).
@@
To test our method on real-world data, we predicted causal genes at 36 genome-wide significant loci for coronary artery disease (CAD) from the CARDIOGRAMC4D meta-analysis study [36] using eQTL data from seven vascular and metabolic tissues from the STARNET study [26] (see Methods for details). Results are presented here for the least squares estimator; the GMM estimates are predominantly consistent and are provided in the Supplementary Information.
@@
Fig 4: Predicted causal genes and their tissue-specific effects on CAD at genome-wide significant CAD loci. (A) Causal effects from univariate MR of genes at 17 CAD GWAS loci with effects on a single gene in a single tissue. (B) Causal effects from tissue-specific, univariate MR of genes at seven CAD GWAS loci with cis-regulatory effects on a single gene in multiple tissues. (C) Tissue color legend for all panels. (D) Causal effects from tissue-specific MVMR with correlated instrumental variable sets of genes at 12 CAD GWAS loci with cis-regulatory effects on multiple genes in multiple tissues. (E) Scatter plot of tissue-specific causal effect estimates from MVMR vs. univariate MR. (F) Scatter plot of tissue-specific causal effect estimates from MVMR vs. CAD regulary trait concordance (RTC) scores. (G) Scatter plot of tissue-specific causal effect estimates from MVMR in STARNET vs. GTEx.
@@
We confirmed the validity of this threshold using seven genome-wide significant CAD loci which had an effect on expression of a single candidate gene in at least two tissues in the STARNET data, that is, loci where we can plausibly hypothesize that the true causal gene, but not the tissue, is known. We used standard, univariate MR on a tissue-by-tissue basis and found that in all seven cases, the candidate gene had a predicted absolute causal effect size greater than 0.1 in at least one tissue (Fig. 4B, Supplementary Information). These genes included CDKN2B, located in one of the most robust genetic markers for type 2 diabetes, CAD, and myocardial infarction [39]. GWAS SNPs in the CDKN2B locus were associated with expression of CDKN2B in aorta (AOR), subcutaneous fat (SF), and blood, with predicted causal effects $c=-0.75$, $-0.55$, and $0.5$, respectively.
We further computed p-values based on the standard errors of the estimated causal effects. As these p-values are based on one-sample estimates for the standard error and may not reflect variability in two-sample summary-data settings (see Methods for details), we conservatively set a Bonferroni threshold of $p<0.05/150=3\cdot 10^{-4}$, which exclude only five gene-tissue combinations with estimated causal effect $|\hat{c}|>0.1$.
@@
Having thus found a reasonable threshold to define causal genes for CAD, we analyzed 12 genome-wide significant CAD GWAS loci which had an effect on expression of multiple nearby genes in one or more tissues. Similar to Porcu et al. [10], we applied MVMR on a tissue-by-tissue basis using all genes with cis-eQTL associations in a locus in a given tissue as potential exposures in a causal diagram (cf. Fig. 1B,E). We note that at only two of the 12 loci, it was possible to reduce the SNPs to a set of independent instruments using the LD threshold ($r^{2}\leq 0.1$) used by Porcu et al. [10], while still retaining at least as many instruments as exposures. Hence an approach using correlated instrument sets is essential.
@@
In total, 35 candidate causal genes with cis-eQTL associations were located in the 12 CAD GWAS loci with pleiotropic gene regulatory effects, of which 24 had a predicted absolute causal effect size greater than 0.1 in at least one tissue (Fig. 4D, Supplementary Information). To analyze these causal genes more systematically, we compared the predicted causal effects from MVMR against the predicted causal effects from standard MR (Fig. 4E) and regulatory trait concordance (RTC) scores [40] computed by Franzén et al. [26] (Fig. 4F), two univariate methods that ignore the pleiotropic gene regulatory effects. We also conducted a replication analysis for loci with cis-associations in matching tissues in both STARNET and GTEx (Fig. 4G, see Methods for details).
@@
A clear example of a CAD GWAS locus where MVMR and MR give contrasting results is a locus centred around chr 6:12,901,441. In STARNET, the locus has cis-associations with the expression of PHACTR1, GFOD1, TBC1D7, RP1-257A7.4, and RP1-257A7.5 in MAM (Fig. 5A,B). In univariate MR analyses these genes have predicted causal effects ranging from 0.15 for PHACTR1 to 0.31 for GFOD1, and a previous integrative genomics analysis lists all of them as candidate causal genes for this locus [41]. After pruning nearly identical SNPs (LD $r^{2}\geq 0.95$), 12 instruments (with mutual LD ranging from $r^{2}=0.17$ to $r^{2}=0.86$) were available for conducting MVMR, resulting in a predicted causal effect of $0.19$ for PHACTR1, with all other effects below the threshold of $0.1$ in absolute value, suggesting that PHACTR1 is the only causal gene in arterial tissue at this locus.
Fig 5: Example CAD GWAS loci with pleitropic gene regulatory effects and predicted causal genes. Each row shows the predicted causal effects from tissue-specific MVMR of genes with cis-eQTL associations in a CAD GWAS locus of interest across seven vascular and metabolic tissues from the STARNET study (left) and the genetic architecture and cis-regulatory pleiotropy of the locus. The heatmaps show the causal effects on the liability on a standardized scale. Only tissues with non-zero cis-eQTL effects in the STARNET study are included. Values identically zero indicate gene-tissue combinations not included as exposures in the model (no cis-eQTL association in that tissue). The LocusZoom plots show for each SNP in the locus the GWAS $-\log_{10}(p-value)$ (y-axis) and associated cis-eQTL genes (symbol, see legends) in a selected tissue. (A, B) Locus centred on chr 6:12,901,44 with LocusZoom plot for MAM. (C, D) Locus centred on chr 15:79,141,784 with LocusZoom plot for AOR. (E, F) Locus centred on chr 2:85,809,989 with LocusZoom plot for AOR.
@@
Another example is the locus centred around chr 15:79,141,784, where functional studies support a causal and proatherogenic role for ADAMTS7 [42]. In STARNET, the locus has cis-associations with the expression of ADAMTS7 in AOR, MAM, and VAF(Fig. 5C,D). In all three tissues, the locus also has cis-associations with the expression of CTSH. After pruning nearly identical SNPs (LD $r^{2}\geq 0.86$), 9, 25, and 10 instruments (with mutual LD ranging from $r^{2}=0.03$ to $r^{2}=0.94$) were available for conducting MVMR in AOR, MAM, and VAF respectively. In the arterial tissues, MVMR confirms that ADAMTS7 is the most likely causal gene (predicted causal effects $0.18$ and $0.14$ vs. $0.025$ and $-0.058$ for CTSH in AOR and MAM, respectively; Fig. 5 C). Replication in GTEx AOR tissue confirmed these results (predicted causal effects $0.18$ and $0.009$ for ADAMTS7 and CTSH, respectively). By contrast, in VAF, both genes are predicted causal with effects in the opposite direction (predicted causal effects $-0.18$ and $0.25$ for ADAMTS7 and CTSH, respectively), the latter due to opposite eQTL effects of the same variants in VAF vs. AOR and MAM. No eQTL associations with either gene were available in GTEx in adipose tissue to confirm these results. The overall picture is further complicated by the fact that the locus also has cis-associations with the expression of CTSH in blood, SKLM, and SF, and with PSM4 in SKLM. The absence of associations with ADAMTS7 in these tissues implies that the association between the locus and CAD is automatically attributed to the other genes (Fig. 5 C). Given the overwhelming functional evidence for ADAMTS7 [42], these results probably reaffirm the importance of having all true exposures included in the model (cf. Fig. 2F-H).
@@
As a final example we consider the locus centred around chr 2:85,809,989. Previous analyses found a candidate causal SNP in this locus located in the promoter region of MAT2A, but since the SNP was associated with expression of multiple genes in this locus, a causal gene could not be predicted [43]. In STARNET, the locus has cis-associations with the expression of MAT2A in AOR and MAM. In both tissues, the locus also has cis-associations with the expression of GGCX, and both genes have previously been listed as candidate causal genes in arterial tissues [41]. After pruning nearly identical SNPs (LD $r^{2}\geq 0.95$), three instruments (with mutual LD ranging from $r^{2}=0.69$ to $r^{2}=0.89$) were available for conducting MVMR. In both AOR and MAM, MVMR suggests MAT2A, and not GGCX is the causal gene (predicted causal effects $-0.23$ and $-0.33$ vs. $0.02$ and $0.16$ for GGCX in AOR and MAM, respectively; Fig. 5 E,F). The overall picture is again less clear, since the locus also has cis-associations with the expression of GGCX and VAMP8 in liver, and with GGCX alone in SF and VAF. Interestingly, in liver, neither gene is predicted causal (predicted causal effects $-0.08$ and $-0.06$; Fig. 5 E), while in the adipose tissues, the entire association between the locus and CAD is automatically attributed to GGCX (predicted causal effects $-0.14$ in both tissues; Fig. 5E).
@@ Discussion
Using the method of path coefficients we showed that the MVMR causal diagram with correlated instruments satisfies the instrumental set condition, a classical result by Brito and Pearl [24] that guarantees the identifiability of the direct causal effects of all exposures in the model on the outcome variable of interest. Moreover, the effects solve a regression of (univariate) regression coefficients, and the only requirement is that the set of instruments consists of (or tags) at least as many causal variants as the number of candidate genes included in the model. We further showed that the standard two-stage least squares estimator of the causal effects in MVMR is part of a generalized method of moments (GMM) family of finite-sample estimators for the theoretical identification equation, as is a least squares estimator that simply replaces the theoretical univariate regression coefficients by their finite sample estimates, results which again do not depend on the presence or absence of instrument correlations.
Extensive simulations confirmed the validity and usefulness of these theoretical results. Most surprisingly perhaps was the finding that the variance in causal effect estimates remained small, even at high correlation values between the instruments. Moreover we saw no benefit in using the computationally more expensive two-stage least squares estimator instead of the simpler least squares estimator – both were unbiased and had identical variance in all our simulations. As expected, horizontal pleiotropy or misspecified LD matrices introduce bias in the estimated causal effects, as in all applications of MR, but using conservative simulation parameters, we observed convergence to the correct estimates at sample sizes at the upper end of what is currently available.
We applied our method to predict causal genes for CAD using eQTL data from seven vascular and metabolic tissues obtained from 600 coronary artery bypass grafting surgery patients in the STARNET study. While predicting CAD risk using expression data from CAD cases can be criticized, we have shown previously that more gene expression traits are associated with CAD risk loci in STARNET cases than in comparable disease-unspecific samples (e.g., GTEx) [26], that eQTLs inferred from CAD cases explain a large proportion of heritability of CAD risk [44, 45], and that variation in eQTL-associated genes and gene networks correlates with the extent of atherosclerosis in CAD cases [46, 45], suggesting that variation in gene expression and disease severity in CAD cases can indeed reflect underlying causal effects on CAD risk. This was further supported by the current finding that causal effects estimated from STARNET and GTEx samples for the same genes and tissues correlated well.
Our analysis of the STARNET data illustrated the importance of being able to apply MVMR with correlated instrument sets. Out of 36 genome-wide significant loci with cis-eQTL associations in STARNET, only 17 were associated with a single gene in a single tissue, 7 with a single gene in multiple tissues, and 12 with multiple genes in multiple tissues. A deeper look into the predictions at some of these loci with widespread regulatory effects showed both the strengths and limitations of applying MVMR in this context.
@@
The fact that regulatory pleiotropy can extend over multiple cell types and tissues implies that MVMR must be run strictly speaking by including all gene-tissue pairs with cis-eQTL associations in a given GWAS locus as exposures in a single, large causal modelin order to limit the risk of horizonal pleiotropy. However, even though instrument sets can be highly correlated and the STARNET data contains only seven tissues, it was almost never possible to run such large models, as the requirement on the number of causal variants included in or tagged by the instrument set still stands. Moreover, data from some potentially relevant cell types or tissues will almost always be missing. To overcome these limitations, we see two ways forward.
First, we should not consider MVMR to be a hypothesis-free method that can be run across all available tissue eQTL data to discover causal genes de novo. Instead prior knowledge of the relevant tissue and plausible candidate genes must be used to limit both the size of the model and the risk that true causal genes and gene-tissue combinations are missing from the model. For instance, integrative genomics analysis pipelines such as the ones developed by Brænne et al. [43] and Hao et al. [41] combine multiple eQTL and GWAS datasets with functional genome annotation and literature search to compile ranked lists of the most likely causal genes and tissues for an outcome of interest such as CAD. Using MVMR as an additional step in such pipelines to resolve causal effects at loci with eQTL effect on prioritized genes in prioritized tissues can provide additional causal evidence, as illustrated by the PHACTR1 and ADAMTS7 examples.
@@
We focused our analysis on genome-wide significant GWAS loci because the use of eQTL information to identify candidate genes is an important part of modern GWAS protocols [50] and because candidate genes at many GWAS loci have been analyzed using orthogonal data, providing some form of validation for our approach. Needless to say, our approach also could be applied to loci passing less stringent thresholds. More generally, the method could be integrated in existing methods to account for linkage disequilibrium and potential pleiotropic effects in transcriptome-wide association studies using statistical fine-mapping approaches [51, 52, 53]. These methods typically use sparsity-inducing constraints to predict outcome traits from a minimal set of most informative genes. Using results from a causal model such as considered here, either directly in an outcome prediction model, or indirectly as part of the sparsity constraints, could increase the probability that the minimal predictive gene set consists of truly causal genes.
Conclusion